# Reading Subtext: Evaluating Large Language Models on Short Story Summarization with Writers

### v1

@@ Abstract
We evaluate recent Large language Models (LLMs) on the challenging task of summarizing short stories, which can be lengthy, and include nuanced subtext or scrambled timelines. Importantly, we work directly with authors to ensure that the stories have not been shared online (and therefore are unseen by the models), and to obtain informed evaluations of summary quality using judgments from the authors themselves. Through quantitative and qualitative analysis grounded in narrative theory, we compare GPT-4, Claude-2.1, and LLama-2-70B. We find that all three models make faithfulness mistakes in over 50% of summaries and struggle to interpret difficult subtext. However, at their best, the models can provide thoughtful thematic analysis of stories. We additionally demonstrate that LLM judgments of summary quality do not match the feedback from the writers.
@@ 1 Introduction
For our narrative form, we use fiction short stories as they present several interesting challenges. Fiction writing does not follow a clear intro-body-conclusion format and instead may follow a nonlinear timeline, hint at things only abstractly, or deliberately mislead the reader. Fiction may use multiple language varieties and creative language. For example, consider this quote from Toni Morrison’s Beloved: “The pieces I am, she gather them and give them back to me in all the right order" Morrison (2004). This sentence uses complex metaphor and African American Language Deas et al. (2023); Grieser (2022) from the late 1800’s to express a beautiful relationship between two characters. Finally, fiction short stories can be longer than an LLM’s context window. Here, we consider stories up to 10,000 tokens long.
@@
Evaluations of narrative summarization on long documents have been scarce due to several key challenges: 1) Narrative text is generally either in the public domain (and therefore likely in LLM training data) or under copyright, and 2) Holistic summary evaluation has been prohibitively difficult due to a lack of reliable automatic metrics Fabbri et al. (2021a); Chang et al. (2023) and complications with human evaluation. For example, it can take someone over an hour to thoroughly read and evaluate just one story and summary, which quickly becomes expensive.
@@
Chang et al. (2023) make progress on these issues by purchasing recent books (which are less likely to be in models’ training data but may still be discussed in the training data), and developing an LLM-based metric for evaluating summary coherence. We take this a step further by working directly with experienced creative writers, and thus are able to: 1) use stories that are not discussed or present in training data, 2) expand beyond evaluation of coherence to attributes focused on narrative understanding like faithfulness and thematic analysis, and 3) use human rather than model judgments.
@@
We evaluate three LLMS – GPT-4, Claude-2.1, and LLama-2-70b – on 25 short stories. We ask authors to evaluate the summaries of their own unpublished stories [^1] We release our code, the writer evaluation responses, and the story/summary errors here: https://github.com/melaniesubbiah/reading-subtext. since they are experts in what they have written, focusing on coherence, faithfulness, coverage of important details, and useful interpretation in the summaries. We then present an analysis of the quantitative and qualitative feedback. Our contributions include:
  1. 1.
    A characterization of LLM narrative understanding based on narrative theory Mandler and Johnson (1977); Piper et al. (2021), resulting in directions for future research.
  2. 2.
    A methodology for evaluating LLM summaries when the lines between test and training data have blurred, as well as a demonstration that LLM judgments cannot replace skilled human evaluators for this task.
  3. 3.
@@
    Informed ratings of faithfulness and the interpretive quality of LLM summaries using experienced creative writers and unseen short stories.
@@ 2 Related Work
Narrative Summarization. Interest in long narrative summarization has steadily grown in the last several years. Ladhak et al. (2020) first introduced the task of summarizing chapters from novels, which was then expanded into the full BookSum dataset Kryściński et al. (2021) and used in early studies of RLHF Wu et al. (2021). Most similar to our work is SQuALITY Wang et al. (2022), which also focuses on short stories. However, their stories are sourced from Project Gutenberg which is likely memorized by LLMs. Across other areas of narrative understanding, recent datasets have been proposed in screenplays Chen et al. (2021), poetry Mahbub et al. (2023), and theory of mind evaluation Xu et al. (2024).
Summarization Evaluation. Evaluating summary quality in any of these areas is a challenge. Work like Fabbri et al. (2021a) has shown the flaws in many of the traditional automatic metrics, prompting a move toward model-based reference-free methods and fine-grained span analysis. LongEval Krishna et al. (2023), QAFactEval Fabbri et al. (2021b), FALTE Goyal et al. (2022a), and FActScore Min et al. (2023) have furthered methods in faithfulness, while work like SNaC Goyal et al. (2022b) and BooookScore Chang et al. (2023) have focused on coherence. Chang et al. (2023) is most similar to our work, but they focus on a GPT prompting strategy for evaluation, evaluate on recently published books, and only evaluate summary coherence. Several studies have benchmarked LLMs on summarization tasks Tang et al. (2024); Zhang et al. (2023a); Tang et al. (2022); Jahan et al. (2024); Liu et al. (2023); Pu et al. (2023), including ones Zhang et al. (2023b); Goyal et al. (2023) that have shown performance is saturated on the commonly used CNN/DM news summarization dataset Hermann et al. (2015).
@@
Chakrabarty et al. (2023a) – 10 writers, 12 stories
Chakrabarty et al. (2023b) – 17 writers, 30 stories
@@
These studies use similar numbers of writers and stories to us given the challenges/cost involved.
@@ 3 Writers and Data
Given the problems with using stories that may have contents or analysis available online, we want to find skilled writers with unpublished stories they are willing to share. We therefore recruit writers through MFA listservs, posts on X, and direct emails. Once writers express interest, we screen for skill by writer education level and/or portfolio. Finally, we obtain their informed consent by sharing an infosheet on the study, which explains the task of sharing their stories and evaluating summaries, and includes how their work will be used as outlined in IRB protocol AAAU8875 (see full infosheet in Appendix A). We ask two questions to determine the background of the writers: 1) Do you have an undergraduate or MFA writing degree?, and 2) Have you previously published any of your writing? To protect writers’ anonymity, we use anonymous IDs to collect and store all the data. In line with Chakrabarty et al. (2023a) and Ippolito et al. (2022), who use 8 and 13 writers respectively, our recruited group consists of 9 skilled writers – 8/9 possess a writing degree, and 7/9 have previously published writing.
Each writer is given the opportunity to submit short stories that they have written and not published anywhere online. We limit the writers to five stories at most, so one writer does not dominate the data. Four writers choose to submit five stories, and five submit one story, resulting in a dataset of 25 stories. We compensate the writers for their submitted stories and evaluations. Table 1 shows some summary statistics for this data. None of the stories exceed 9900 tokens in length, and for the purpose of analysis, we consider performance across three length buckets: short (<3300 tokens), medium (3300-6600 tokens), and long (6600-9900 tokens). Token count is computed as the average of the GPT-4 tokenizer [^2] https://cookbook.openai.com/examples/how_to_count_tokens_with_tiktoken count and the Llama2 tokenizer [^3] https://huggingface.co/docs/transformers/model_doc/llama2#transformers.LlamaTokenizer count as Claude does not have a publicly available tokenizer.
@@ 4 Summary Generation
Figure 1: The two different methods used for summarization and the associated prompts for the models. GPT-4 and Claude have sufficient input context to summarize a whole story, whereas Llama has to use a chunk-then-summarize approach for longer stories.
@@
Model Coverage Density Compression
GPT-4 72.52% 1.33 7.45
Claude 73.65% 1.26 10.86
Llama 66.09% 1.08 7.66
Table 2: We report the average coverage, density, and compression scores for the summaries generated by each model in relation to the stories.
@@
Protecting the writers’ unpublished work is paramount when using each of these three models. We do not want their stories to be saved or trained on by OpenAI, Anthropic, or HuggingFace. We inform the writers of this risk in the consent form and complete the available forms/settings with each of these companies to request our data not be stored for long periods or used for training.
For each story, we then generate three summaries – one from each of the three models. For GPT-4 and Claude, we use one prompt (see Figure 1) as they both have a context length long enough to process a whole short story. Claude refuses to summarize two of the stories which do not meet its content restrictions [^5] For all of the Claude results, we remove the values for these two summaries to show representative numbers for the summaries it did produce. We also remove one coherence score which was inaccurate due to a bug in the annotation interface.. One of these stories is about someone being shot and the other involves sex and robbery.
@@
Table 1 displays summary statistics for the generated summaries. Notice that our prompt asks models to use about 400 words and Claude undershoots but comes closest to this target while GPT-4 and Llama are 100 words over on average. See Appendix B for more prompting details and experiment costs.
We also compute the average coverage, density, and compression metrics for each model’s summaries (see Table 2). Coverage is the percent of words in the summary that come from the story, density is the average length of segments that directly match the story text, and compression is how many times shorter the summary is in relation to the story Fabbri et al. (2021a); Grusky et al. (2018). These metrics show that Llama is the least extractive and Claude is the most effective at compressing the information.
5 Human Evaluation
We compare models using a combination of quantitative and qualitative evaluation.
5.1 Quantitative
We assess quality in terms of four attributes:
@@ 5.1 Quantitative
Coverage, faithfulness, and coherenece are commonly evaluated in summarization (e.g., Zhang et al. (2023b)), and we add analysis which is important for capturing narrative. Each writer is shown the summaries of their own stories to evaluate as they are deeply familiar with the contents and can therefore judge aspects like faithfulness and analysis quickly and accurately. The writers rate each of these attributes on a Likert scale from 1 to 4 with some guidance on what each score means (see Figure 2, full interface shown in Appendix D).
For each story/summary pair, we also ask GPT-4 and Claude (Llama’s context window is too short) these four questions with the same wording to see how their answers compare to the opinions of the writers (see Appendix B for prompts). We additionally run BooookScore as another comparison point for LLM-based coherence evaluation Chang et al. (2023).
@@
Once each writer has read and evaluated each summary for a story, we ask them to rank the three summaries from the three different models in order of their preference.
Lastly, we analyze the complexity of the story wording using the Flesch-Kincaid readability test Kincaid et al. (1975). This score estimates the number of years of formal education (grades 1 through 12) one might need to understand the writing easily. It is based on average word and sentence length, so while it captures something about the complexity of the wording, it is in no way a measure of the overall quality of the writing. We explore whether the reading-level of the writing affects the summary quality (see examples of writing with assigned grade levels in Figure 3).
@@
5.2 Qualitative
In addition to the ratings, the writers are also asked to provide open-ended feedback on each summary. We read all of the stories/summaries in full and assemble a list of paired text from the stories and summaries that support the writers’ lower scores and written feedback.
To examine which aspects of story content are most often missed in summaries, we ask annotators to categorize the instances of faithfulness errors from this list using elements from narrative theory for evaluating narrative understanding in children Xu et al. (2022); Paris and Paris (2003); Mandler and Johnson (1977). We define the categories for the errors as follows. An error introduces a detail that would incorrectly answer one of the narrative understanding questions in these categories:
Feeling – Emotions/reactions/internal state of characters. Answers a question like, “How did X react?” or “What was X thinking?”
Causation – Causal relationship of events. Answers a question like, “Why did Y happen?” or “What was the result of Y?”
Action – Behavior of a character. Answers a question like, “What did X do?”
Character – Identity or nature of a character. Answers a question like, “How would you describe character X?” or “Who is X?”
Setting – Story world and the time/place of events. Answers a question like, “Where/when did X happen?” or “What is the setting of the story?”
We sample 20 faithfulness errors for each model from across all its summaries. For each error, we ask three Natural Language Processing graduate students to first label whether they agree the details described constitute a faithfulness error, and if so, to categorize it (see full instructions in Appendix D). The Fleiss Kappa score is .51.51.51.51 indicating moderate inter-annotator agreement. For analysis, we use the majority decision for each error, and label it as no-consensus if multiple annotators thought it was not a faithfulness error or all three annotators gave it a different label. We also ask GPT-4 and Claude the same questions using the same prompting method as for answering the quantitative questions but with the instructions for this task (see Appendix B for prompts).
To examine how writing style affects model summarization, we use Genette’s model of narrative elements Piper et al. (2021); Genette (1980):

### v3

@@ Abstract
We evaluate recent Large Language Models (LLMs) on the challenging task of summarizing short stories, which can be lengthy, and include nuanced subtext or scrambled timelines. Importantly, we work directly with authors to ensure that the stories have not been shared online (and therefore are unseen by the models), and to obtain informed evaluations of summary quality using judgments from the authors themselves. Through quantitative and qualitative analysis grounded in narrative theory, we compare GPT-4, Claude-2.1, and LLama-2-70B. We find that all three models make faithfulness mistakes in over 50% of summaries and struggle with specificity and interpretation of difficult subtext. We additionally demonstrate that LLM ratings and other automatic metrics for summary quality do not correlate well with the quality ratings from the writers.
@@ 1 Introduction
For our narrative form, we use fiction short stories as they present several interesting challenges. Fiction writing does not follow a clear intro-body-conclusion format and instead may follow a nonlinear timeline, hint at things only abstractly, or deliberately mislead the reader. Fiction may use multiple language varieties and creative language. For example, consider this quote from Toni Morrison’s Beloved: “The pieces I am, she gather them and give them back to me in all the right order" Morrison (2004). This sentence uses complex metaphor and African American Language Deas et al. (2023); Grieser (2022) from the late 1800’s to express a beautiful relationship between two characters. Finally, fiction short stories can be longer than an LLM’s context window. Here, we consider stories up to 10,000 tokens long, which fit within the context window of the paid LLMs we use but are too long for the context window of the open-source LLM we use.
@@
Evaluations of narrative summarization on long documents have been scarce due to several key challenges: 1) Narrative text is generally either in the public domain (and therefore likely in LLM training data) or under copyright, and 2) Holistic summary evaluation has been prohibitively difficult due to a lack of reliable automatic metrics Fabbri et al. (2021); Chang et al. (2024) and complications with human evaluation. For example, it can take someone over an hour to thoroughly read and evaluate just one story and summary, which quickly becomes expensive.
@@
Chang et al. (2024) make progress on these issues by purchasing recent books (which are less likely to be in models’ training data but may still be discussed in the training data), and developing an LLM-based metric for evaluating summary coherence. We take this a step further by working directly with experienced creative writers, and thus are able to: 1) use stories that are not discussed or present in training data, 2) expand beyond evaluation of coherence to aspects of narrative understanding like faithfulness and thematic analysis, and 3) use human rather than model judgments.
@@
We evaluate three LLMS – GPT-4, Claude-2.1, and LLama-2-70B – on 25 short stories. We ask authors to evaluate the summaries of their own unpublished stories [^1] We release our code, the writer evaluation responses, and the story/summary errors here: https://github.com/melaniesubbiah/reading-subtext since they are experts in what they have written, focusing on coherence, faithfulness, coverage of important details, and useful interpretation in the summaries. We then present an analysis of their evaluation.
The key features of our work are:
1.)
Span-level, summary-level, and story-level evaluation of LLM summaries of short stories.
@@
2.) Experienced creative writers as evaluators and short stories unseen by LLMs.
3.) Exploration of LLM ability to analyze and interpret narrative.
Our key findings are:
1.)
GPT-4 and Claude can produce excellent summaries, but only about half the time.
2.) LLMs struggle with specificity, interpreting subtext, and unreliable narrators.
3.) LLM judgments cannot replace skilled human evaluators for this task.
@@ 2 Related Work
Narrative Summarization. Interest in long narrative summarization has steadily grown in the last several years. Ladhak et al. (2020) first introduced the task of summarizing chapters from novels, which was then expanded into the full BookSum dataset Kryscinski et al. (2022) and used in early studies of RLHF Wu et al. (2021). Most similar to our work is SQuALITY Wang et al. (2022), which also focuses on short stories. However, their stories are sourced from Project Gutenberg which is likely memorized by LLMs. Across other areas of narrative understanding, recent datasets have been proposed in screenplays Chen et al. (2022), poetry Mahbub et al. (2023), and theory of mind evaluation Xu et al. (2024).
Summarization Evaluation. Evaluating summary quality in any of these areas is a challenge. Work such as Fabbri et al. (2021) has shown the flaws in many of the traditional automatic metrics, prompting a move toward model-based reference-free methods and fine-grained span analysis. LongEval Krishna et al. (2023), QAFactEval Fabbri et al. (2022), FALTE Goyal et al. (2022a), and FActScore Min et al. (2023) have furthered methods in faithfulness, while work like SNaC Goyal et al. (2022c) and BooookScore Chang et al. (2024) have focused on coherence. Chang et al. (2024) is most similar to our work, but they focus on a GPT prompting strategy for evaluation, evaluate on recently published books, and only evaluate summary coherence. Several studies have benchmarked LLMs on summarization tasks Tang et al. (2024b); Zhang et al. (2023); Tang et al. (2023); Jahan et al. (2024); Liu et al. (2024); Pu et al. (2023), including ones Zhang et al. (2024); Goyal et al. (2022b) that have shown performance is saturated on the commonly used CNN/DM news summarization dataset Hermann et al. (2015).
@@
Chakrabarty et al. (2024) – 10 writers, 12 stories
Chakrabarty et al. (2023) – 17 writers, 30 stories
@@
These studies use similar numbers of writers and stories as us given the challenges/cost involved.
@@ 3 Writers and Data
To avoid using stories that may have contents or analysis available online, we find skilled writers with unpublished stories they are willing to share. We recruit writers through MFA listservs, posts on X, and direct emails. Once writers express interest, we screen for skill by writer education level and/or portfolio. Finally, we obtain their informed consent by sharing an infosheet on the study, which explains the task of sharing their stories and evaluating summaries, and includes how their work will be used as outlined in IRB protocol AAAU8875 (see full infosheet in Appendix A). We ask two questions to determine the background of the writers: 1) Do you have an undergraduate or MFA writing degree?, and 2) Have you previously published any of your writing? To protect writers’ anonymity, we use anonymous IDs to collect and store all the data. In line with Chakrabarty et al. (2024) and Ippolito et al. (2022), who use 8 and 13 writers respectively, our recruited group consists of 9 skilled writers – 8/9 possess a writing degree, and 7/9 have previously published writing.
Each writer is given the opportunity to submit short stories that they have written and not published anywhere online (see interface in Appendix C). We limit the writers to five stories at most, so one writer does not dominate the data. Four writers choose to submit five stories, and five submit one story, resulting in a dataset of 25 stories. We compensate the writers for their submitted stories and evaluations. Table 1 shows some summary statistics for this data. None of the stories exceed 9900 tokens in length, and we show statistics by different story length buckets: short (<3300 tokens), medium (3300-6600 tokens), and long (6600-9900 tokens). Token count is computed for each story as an average of the GPT-4 tokenizer [^2] https://cookbook.openai.com/examples/how_to_count_tokens_with_tiktoken count and the Llama2 tokenizer [^3] https://huggingface.co/docs/transformers/model_doc/llama2#transformers.LlamaTokenizer count. Claude does not have a publicly available tokenizer, so we cannot include its token count in this average.
We also explore how similar the stories may be to the LLM training data by measuring their perrplexity. We cannot compute perplexity using the GPT-4 and Claude APIs, so we get an approximate number by using Llama-2-7B (Float16), which we can run on our own compute. This model assigns perplexity scores between 9.2 and 21.6 to the stories. For reference, the first chapter of Pride and Prejudice (a story in Llama-2’s training data) has a perplexity score of 1.3, which is just under the model’s training loss. These scores therefore validate that the stories are unfamiliar to the models.
@@ 4 Summary Generation
Figure 1: The two different methods we use for summarization and the associated prompts for the models. GPT-4 and Claude have sufficient input context to summarize a whole story, whereas Llama has to use a chunk-then-summarize approach for longer stories.
Model Coverage Density Compression 2-grams 3-grams 4-grams 5-grams LCS
GPT-4 72.52% 1.33 7.45 22.53% 5.18% 1.76% 0.71% 5.72
Claude 73.65% 1.26 10.86 19.80% 3.96% 1.25% 0.50% 4.16
Llama 66.09% 1.08 7.66 16.33% 3.04% 0.96% 0.32% 4.40
Table 2: We report a variety of metrics for the summaries in relation to the stories: coverage, density, compression, percent of n-gram overlap, and word count of the longest common substring. Each metric is averaged across the summaries generated by each model.
@@
Protecting the writers’ unpublished work is paramount when using each of these three models. We do not want their stories to be saved or trained on by OpenAI, Anthropic, or HuggingFace. We inform the writers of this risk in the consent form (see Appendix A) and complete the available forms/settings with each of these companies to request our data not be stored for long periods or used for training.
For each story, we then generate three summaries – one from each of the three models. For GPT-4 and Claude, we use one prompt (see Figure 1) as they both have a context length long enough to process a whole short story. Claude refuses to summarize two of the stories which do not meet its content restrictions [^5] For all of the Claude results, we remove the values for these two summaries to show representative numbers for the summaries it did produce. We also remove one coherence score which was inaccurate due to a bug in the interface.. One of these stories is about a shooting and the other involves sex and robbery.
@@
Table 1 displays summary statistics for the generated summaries. Notice that our prompt asks models to use about 400 words and Claude undershoots but comes closest to this target while GPT-4 and Llama are 100 words over on average. See Appendix B for full prompting details and costs.
In Table 2, we report the average coverage, density, and compression metrics for each model’s summaries. Coverage is the percent of words in the summary that come from the story, density is the average length of segments that directly match the story text, and compression is how many times shorter the summary is in relation to the story Fabbri et al. (2021); Grusky et al. (2018). These metrics show that Llama is the least extractive and Claude compresses the information the most.
In addition to these commonly used metrics, we further investigate how much the models copy directly from the stories in Table 2. We report the average percent of n-grams in the summaries that match n-grams in the stories, and the average word count of the longest substring that exactly matches wording from the story. For the n-gram matching, we remove punctuation and lower-case and stem words. We see that GPT-4 copies a significant amount of wording from the stories. On average, it produces summaries with almost 6-word long exact-match substrings and it has the highest percentages of n-gram overlap. These numbers indicate that models are not copying long quotes, but they are copying a non-trivial amount of unique phrasing. None of the summaries use quotation marks to appropriately attribute copied text.
5 Evaluation
We compare models using a combination of span-level, summary-level, and story-level evaluation. We assess summary quality in terms of four attributes:
@@ 5 Evaluation
Coverage, faithfulness, and coherence are commonly evaluated in summarization (e.g., Zhang et al. (2024)), and we add analysis, which is important for capturing narrative. Each writer is shown the summaries of their own stories to evaluate as they are deeply familiar with the contents and can therefore judge aspects like faithfulness and analysis quickly and accurately.
5.1 Span-Level Error Categorization
We ask the writers to highlight spans in the summaries they view as errors and categorize them (see Appendix D for interface and cost). We do not make this task mandatory as it requires significantly more time from the writers than the summary-level evaluation discussed next. 7/9 writers choose to participate, which results in 69/75 summaries with span-level annotations.
For faithfulness, we use error categories inspired by elements from narrative theory for evaluating narrative understanding in children Xu et al. (2022); Paris and Paris (2003); Mandler and Johnson (1977). We determine these categories are well-defined as in prior work by asking three NLP researchers to categorize a sample of 60 faithfulness errors. These annotators achieve moderate inter-annotator agreement for the labels (Fleiss-Kappa of .51), indicating the categories are valid. For coherence, we use categories defined and validated by Chang et al. (2024). For coverage, we cannot highlight spans that should have been included, so we can only focus on things that should not have been included or are non-specific. The full list of error categories are defined as:
Coverage
Insignificant - Does not need to be included and makes the summary less readable
Vague - Important but covered in a vague way
Faithfulness
Feeling - Inaccurate about a character’s emotions/reaction/internal state, incorrectly answers a question like "How did X react" or "What was X thinking?"
Character - Inaccurate about the identity or nature of a character, incorrectly answers a question like "How would you describe X?" or "Who is X?"
Causation - Inaccurate about the causal relationship of events, incorrectly answers a question like "Why did Y happen?" or "What was the result of Y?"
Action - Inaccurate about the behavior of a character, incorrectly answers a question like "What did X do?"
Setting - Inaccurate about the details of the story world and the time/place of events, incorrectly answers a question like "Where/when did X happen?" or "What is the setting of the story?"
Coherence
Inconsistent - Inconsistent with other details in the summary
Abrupt Transition - Transitions suddenly to a new scene without a relevant connection
Missing Context - Introduces a new character, event, or object without enough context/detail to understand it
Repetition - Unnecessary repetition of detail
Analysis
Unsupported - Interprets the story but the conclusions do not make sense with or are unsupported by the story
5.2 Summary-Level Ratings
Rating Cover. Faithful. Coher. Analys.
1 1.00 3.33 - 0.00
2 1.92 1.32 1.00 2.33
3 1.48 0.87 1.36 1.50
4 0.50 0.89 0.76 1.22
Table 3: Average number of span-level errors by attribute labeled in summaries given each Likert score rating for that attribute.
We ask the writers to rate each of the four attributes on a Likert scale from 1 to 4 with some guidance on what each score means (see Figure 2, full interface and cost shown in Appendix C). This step is completed before the span-level annotation to keep conditions equal between writers who opt in or out of the span-level annotation.
To help interpret the Likert scores, in Table 3, we present a breakdown of the average number of span-level errors annotated in each attribute when a summary is given a certain Likert score rating for that attribute. We can see that for most attributes, there is an increase in average error count for an attribute as the rating assigned to that attribute decreases. The number of 1-ratings is quite small for many attributes so the relationship between number of errors and rating is noisier for that bucket. Overall, these numbers support our 4-point Likert scale and demonstrate that the summary-level ratings correspond to meaningful differences in the number of errors in a summary.
@@ 5.2 Summary-Level Ratings
Once each writer has read and evaluated each summary for a story, we ask them to rank the three summaries from the three different models in order of their preference. In addition to the ratings and ranking, the writers provide open-ended feedback on each summary, which we include quotes from throughout our discussion of the results.
We also explore automatic metrics for rating summaries. We try simple automatic metrics, using ROUGE Lin (2004) and BERTScore Zhang et al. (2020) against the full story as a reference. We try recent attribute-specific metrics designed for faithfulness: AlignScore Zha et al. (2023), UniEval Zhong et al. (2022), and MiniCheck Tang et al. (2024a). Finally, we ask GPT-4 and Claude (Llama’s context window is too short) the four questions shown in Figure 2 with the same wording to see how their answers compare to the opinions of the writers (see Appendix B for prompts). We run BooookScore and FABLES as another comparison point for LLM-based coherence and faithfulness evaluation respectively Chang et al. (2024); Kim et al. (2024). For FABLES, we average the labels for all the claims in a summary to get a score for the whole summary.
5.3 Story-Level Style Effects
@@
We examine how writing style at the story level affects model summarization to see if certain types of stories are harder for models to summarize well. First, we use Genette’s model of narrative elements Piper et al. (2021); Genette (1980):
@@
Finally, we analyze the complexity of the story wording using the Flesch-Kincaid readability test Kincaid et al. (1975). This score estimates the number of years of formal education (grades 1 through 12) one might need to understand the writing easily. It is based on average word and sentence length, so while it captures something about the complexity of the wording, it is in no way a measure of the overall quality of the writing. We explore whether the reading-level of the writing affects the summary quality (see examples of writing with assigned grade levels in Figure 3).